suppose a ray divides 112 degree angle into two angles where one is 22 less than 5 times the other. what are the measures of the two angles?
Answer by cleomenius(959)

he two angles together will add to 112 degrees; so we solve for x.
x - 22 + 5x = 112
6x = 134
x = 22.33 The smaller angle.
(5)22.33 = 111.65 - 22 = 89.65 The larger angle
89.65 + 22.33 = 111.98 (round off error).
This does check.
Cleomenius.